OpenStudy (anonymous):

Find all of the zeros of the equations below. Show all of your work. 1) x3 – 5x2 + 6x = 0 2) x4 – 8x2 – 9 = 0 3) 5x3 – 5x = 0 This is the only thing I have left to do and I have no idea how to do it, please help

OpenStudy (anonymous):

1.\[x \left( x^2-5x+6 \right)=0,either~ x=0,or~x ^{2}-3x-2x+6=0,make factors\]

OpenStudy (anonymous):

Can you factor these equations?

OpenStudy (anonymous):

if you factor them you get: for number one you get: \[x(x^2-5x+6)=0\] using the zero factor property you know that either x = 0 or \[x^2-5x+6=0\] so so far we know that 0 is a solution now solve the quadratic: \[(x-3)(x-2)=0\] x=3 or x=2 So the zeroes for the first equation are 0,2, and 3

OpenStudy (anonymous):

\[x^4-8x^2-9=0\] \[put ~x^2=y,x^4=y^2\] \[y^2-8y-9=0\] 1*-9=-9 9-1=8 9*-1=-9 -8y=-(9-1)y=-9y+1y \[y^2-9y+1y-9=0,y \left( y-9) \right)+1\left( y-9 \right)=0\] \[\left( y-9 \right)\left( y+1 \right)=0\] either y-9=0,y=9 \[x ^{2}=9,x=\pm 3\] or y+1=0,y=-1 \[or~x ^{2}=\iota ^{2},x=\pm \iota \]

OpenStudy (anonymous):

3. \[5x ^{3}-5x=0,5x \left( x ^{2}-1 \right)=0,5x \left( x+1 \right)\left( x-1 \right)=0\] Hence x=0,-1,1